OpenStudy (anonymous):

I need help with this. Can someone help me please. A wheel makes 6 7/15 revolutions per minute. If it rotates for 80 minutes, how many revolutions does it make?

OpenStudy (anonymous):

67/15 x 80 = 357.33 revolutions

ganeshie8 (ganeshie8):

since per minute its making 6 full revolutions + 7/15th of a revolution, in 80 minutes it would make 80*6 + 80*7/15 revolutions

OpenStudy (anonymous):

How you multiply 7/15

OpenStudy (anonymous):

Im still confuse with this problem.

ganeshie8 (ganeshie8):

which part u confuse wid ? did u get how it makes ( 80*6 + 80*7/15 ) revolutions in 80 minutes ?

OpenStudy (anonymous):

i kind of got it but not really. Algebra is very complicated

ganeshie8 (ganeshie8):

if we rush, it looks complicated. go slow, then it would make sense

OpenStudy (anonymous):

80*6=480+80=560*7/15 im comfuse with the 7/15

ganeshie8 (ganeshie8):

ok, so u get how its ( 80*6 + 80*7/15 ) revolutions in 80 minutes, but u want to knw how to simplify it, right ?

OpenStudy (anonymous):

yes, your right. how too simplify it

ganeshie8 (ganeshie8):

80*6 + 80*7/15 480 + 560/15

ganeshie8 (ganeshie8):

we can simplify further.

OpenStudy (anonymous):

How would you answer the problem.

ganeshie8 (ganeshie8):

80*6 + 80*7/15 480 + 560/15 480 + (555 + 5)/15

ganeshie8 (ganeshie8):

80*6 + 80*7/15 480 + 560/15 480 + (555 + 5)/15 480 + 555/15 + 5/15

ganeshie8 (ganeshie8):

fine, so far ?

ganeshie8 (ganeshie8):

i did nothing fancy, i just wrote 560 = 555 + 5

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

is the answer 560/15

ganeshie8 (ganeshie8):

15 goes in 555 , right ?

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

80*6 + 80*7/15 480 + 560/15 480 + (555 + 5)/15 480 + 555/15 + 5/15 480 + 37 + 5/15

ganeshie8 (ganeshie8):

15 goes in 555, 37 times

ganeshie8 (ganeshie8):

80*6 + 80*7/15 480 + 560/15 480 + (555 + 5)/15 480 + 555/15 + 5/15 480 + 37 + 5/15 517 + 5/15

ganeshie8 (ganeshie8):

80*6 + 80*7/15 480 + 560/15 480 + (555 + 5)/15 480 + 555/15 + 5/15 480 + 37 + 5/15 517 + 5/15 517 + 1/3 517 1/3

ganeshie8 (ganeshie8):

we are done wid simplifying !

ganeshie8 (ganeshie8):

see if that makes sense :)

OpenStudy (anonymous):

ok it does make since. I get it lol

ganeshie8 (ganeshie8):

glad to hear friend :)

OpenStudy (anonymous):

thank you so much

ganeshie8 (ganeshie8):

np.. yw !